A company just issued a 10-year bond at a coupon rate of 5.00% payable annually. These securities trade in the market at a yield to maturity of 4.00%. What is the current price of your bonds?
Answer #1

Annual coupon=1000*5%=50

Hence current price=Annual coupon*Present value of annuity factor(4%,10)+$1000*Present value of discounting factor(4%,10)

=50*8.11089578+1000*0.675564169

=$1081.11(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=50[1-(1.04)^-10]/0.04

=50*8.11089578

2.Present value of discounting factor=1000/1.04^10

=1000*0.675564169